The points P(ap2 , 2ap) and Q(aq2 , 2aq) lie on the parabola y2= 4ax. Prove that if PQ is a focal chord then the tangents to the curve at P and Q intersect at right angles at a point on the directrix.


1
Expert's answer
2021-09-20T16:54:15-0400

Solution:

Let's draw a figure as per given data:



Because (P,Q) is focal hord we have:

"{a\cdot p^2-a\cdot q^2 \over 2\cdot a\cdot p-2\cdot a\cdot q)}=""{a\cdot p^2-a\over 2\cdot a\cdot p};"

"p+q=p-{1\over p}" ;

"p\cdot q=-1"

Let us find the intersection of Z(-a,u(p)) tangent at P with asymptota x=-a:

"4\cdot a\cdot x-y^2=0 - equation \space of parabola"

"2\cdot a\cdot dx-y\cdot dy=0;\space \\
dx=x(P)-x(Z)=a\cdot p^2+a;\\
dy=y(P)-y(Z)=2\cdot a\cdot p-u(p);\\
2\cdot a\cdot(a\cdot p^2+a)-2\cdot p\cdot a\cdot (2\cdot a\cdot p-u(p);
\\a\cdot p^2+a-2\cdot a\cdot p^2+p\cdot u(p)=0;\\
u(p)=\frac{a\cdot p^2-a}{p}=a\cdot p-a\cdot \frac{1}{p}=a\cdot (p+q);"

Similarly u(q)=a(q+p)

Thus u(p)=u(q) therefore tangents at P and Q intersecs at Z(-a,a(p+q));

k(P)=dy(p)/dx(P)="{2\cdot a\cdot p-a\cdot (p-\frac{1}{p})\over a\cdot p^2+a }=\frac{1}{p}" is the slope of tangent at P;

Similarly k(Q)="\frac{1}{q}" .

Therefore k(P)"\cdot" k(Q)=-1 and tangents at P and at Q are ortogonal in Z.

Hence, proved.